how do i solve this: 2xy+9x-5

Answer by ReadingBoosters(3246)

Starting with
2xy + 9x - 5
Assuming you mean to add = 0
2xy + 9x - 5 = 0
Isolate y
2xy = -9x + 5
Divide both sides by 2x
y = -9x/2x + 5/2x
Simplify and Rearrange to follow y=mx+b format
y = 5/2x - 9/2
..............................
Delighted to help.
-Reading Boosters
Wanting for others what we want for ourselves.